Two spherical objects have masses of 1.2 x 10^4 kg and 3.6 x 10^3 kg. A

distance of 4500 m separates their centers. Find the gravitational

attraction between them.


1
Expert's answer
2022-02-28T13:40:59-0500

The universal law of gravitation says


"F=G\frac{m_1m_2}{r^2}\\
=6.67*10^{-11}*\frac{1.2*10^{4}*3.6*10^{3}}{4500^2}=1.42*10^{-10}\;\rm N"